Quinine occurs naturally in the bark of the cinchona tree. For centuries it was the only treatment for malaria. Quinine contains two weakly basic nitrogen atoms, with Kb1 = 3.3×10-6 and Kb2 = 1.4×10-9 at 25°C. Calculate the pH of a 0.0136 M solution of quinine in water.
Answer #1

concentration of base B = 0.0136 M

    B   +    H2O     ---------------> BH+   +   OH-

0.0136                                         0             0

0.0136-x                                      x             x

Kb = x^2 / 0.0136 - x

3.3 x 10^-6 = x^2 / 0.0136 - x

x = 2.118 x 10^-4

[OH-] = 2.118 x 10^-4 M

pOH = -log (2.118 x 10^-4 )

        = 3.67

pH = 10.33